A coil of annealed copper wire has 820 loops, the average length of which is 22.86 cm. If the diameter of the wire is 32 mils, calculate the total resistance of the coil at 20 degree celsius.
1
Expert's answer
2019-09-30T09:48:30-0400
"R=N\rho \frac{l}{A}=4N\rho \frac{l}{\pi d^2}"

"R=4(820)(1.72\cdot 10^{-8})\frac{0.2286}{\pi (0.032)^2}=0.004\Omega=4m\Omega"